OpenStudy (anonymous):

Prove this identity, please with steps: sin5x=sinx(cos^2 2x-sin^2 2x) + 2cosxcos2xsin2x

OpenStudy (callisto):

It's easier to start from the right side, i.e. sinx(cos^2 2x-sin^2 2x) + 2cosxcos2xsin2x. Now, consider 2cosxcos2xsin2x, 2cosxcos2xsin2x = cosx(2cos2xsin2x) Can you use the identity \(\sin(2x) = 2\sin(x)\cos(x)\) to simplify the expression?

OpenStudy (anonymous):

cosx(sin(4x))?

OpenStudy (callisto):

Yes, so now, we have \(\sin x(\cos^2( 2x)-\sin^2( 2x)) + \cos(x) \sin(4x)\) For the first term, \(\sin x(\cos^2(2x)-\sin^2(2x))\) Can you use the identity \(\cos(2x) = \cos^2(x) -\sin^2(x)\) to simplify the expression?

OpenStudy (anonymous):

sinx(cos^2 2x - sin^2 2x) = sinx(cos(4x))

OpenStudy (anonymous):

LS= sinxcos4x + cosxsin4x

OpenStudy (anonymous):

RS= sinx(cos(4x)) + cosx(sin4x))

OpenStudy (callisto):

RS= sinx cos(4x) + cosx sin(4x) is right. Now, use the identity \(\sin(a+b)=\sin(a)\cos(b) + \cos(a)\sin(b)\) to simplify the right side, what do you get?

OpenStudy (anonymous):

sin(5x)

OpenStudy (anonymous):

Yes! Proven! THanks so much!

OpenStudy (callisto):

You're welcome! :D